The summer you have designed has four input lines with R1=1, R2=1/2, R3=1/4, R4=1/8and Rf=1, all values given in kΩ. The input lines are set at 0 or 1v. Find V0 in terms of the following set of input values V4=1, V3=0, V2=0 and V1=1


1
Expert's answer
2021-05-31T06:39:45-0400

Voltage at non-inverting input

"\frac{0-V_1}{R_1}+\frac{0-V_2}{R_2}+\frac{0-V_3}{R_3}+\frac{0-V_4}{R_4}+\frac{0-V_0}{R_f}=0"

Making V0 the subject of the formula

"V_0= \frac{-R_f}{R_1}V_1+\frac{-R_f}{R_2}V_2+\frac{-R_f}{R_3}V_3+\frac{-R_f}{R_4}V_4"

Substituting the values

"V_0= \frac{-1}{1}*1+\frac{-1}{0.5}*0+\frac{-1}{1/4}*0+\frac{-1}{1/8}*1"

"V_0= -1-8"

"V_0= -9 V"